# VQ-HPS: Human Pose and Shape Estimation in a Vector-Quantized Latent Space

### v1

@@ Abstract
Human Pose and Shape Estimation (HPSE) from RGB images can be broadly categorized into two main groups: parametric and non-parametric approaches. Parametric techniques leverage a low-dimensional statistical body model for realistic results, whereas recent non-parametric methods achieve higher precision by directly regressing the 3D coordinates of the human body. Despite their strengths, both approaches face limitations: the parameters of statistical body models pose challenges as regression targets, and predicting 3D coordinates introduces computational complexities and issues related to smoothness.
In this work, we take a novel approach to address the HPSE problem. We introduce a unique method involving a low-dimensional discrete latent representation of the human mesh, framing HPSE as a classification task. Instead of predicting body model parameters or 3D vertex coordinates, our focus is on forecasting the proposed discrete latent representation, which can be decoded into a registered human mesh. This innovative paradigm offers two key advantages: firstly, predicting a low-dimensional discrete representation confines our predictions to the space of anthropomorphic poses and shapes; secondly, by framing the problem as a classification task, we can harness the discriminative power inherent in neural networks.
Our proposed model, VQ-HPS, a transformer-based architecture, forecasts the discrete latent representation of the mesh, trained through minimizing a cross-entropy loss. Our results demonstrate that VQ-HPS outperforms the current state-of-the-art non-parametric approaches while yielding results as realistic as those produced by parametric methods. This highlights the significant potential of the classification approach for HPSE.
Figure 1: VQ-HPS formulates the human pose and shape estimation problem as a classification task in a vector-quantized latent space. We present the results of VQ-HPS on two challenging scenarios with in-the-wild conditions and poor illumination, comparing its performance to that of HMR [25], Cliff [36] and FastMETRO [7].
@@ 1 Introduction
In order to alleviate these issues, several works switched to methods inspired by 3D pose estimation models that predict 3D coordinates directly. Earlier methods predicted the 6890 vertices of the full SMPL [40] mesh using graph convolutional neural networks (GCNNs) modeling the mesh structure and focusing on local interaction between neighboring vertices [9, 32]. While [38] used Transformers [62] to model global interactions between joints and vertices, others argued that a hybrid architecture mixing GCNNs and Transformers would enable modeling both local and global interactions [39]. More recently, FastMETRO [7] proposed a transformer-based encoder-decoder architecture to disentangle image and mesh features and predicted 3D coordinates of body joints and a coarse mesh that can be upsampled to the full SMPL body mesh. Significantly different from prior works, LVD [12] proposed an optimization-based approach estimating each vertex position independently by predicting vertex displacement with neural fields. Despite proposing alternatives to model-based approaches, these methods also present some drawbacks: 1) Approaches regressing all vertices of the body mesh at once lack global interaction modeling when using GCNNs [39] and have a very high computational cost when using Transformers [7, 14]; 2) Regression-based methods sometimes output noisy meshes, some of them regress the SMPL parameters from the predicted mesh to obtain smoother predictions, but it comes with a loss of accuracy [7, 32, 9]; 3) LVD [12] is real-time and obtains state-of-the-art results for shape estimation, but is not adapted to extreme poses.
@@
This work introduces a method significantly different from all prior human pose and shape estimation approaches. Instead of predicting the parameters of a human body model or 3D coordinates, we learn to predict a discrete latent representation of 3D meshes, transforming the HPSE into a classification problem, in which we can exploit the discriminative power of transformers. Concretely, for learning our discrete latent representation of meshes, we build on the vector quantized-variational autoencoder (VQ-VAE) [60] framework and adapt it to the fully convolutional mesh autoencoder proposed in [78]. The encoder of the proposed model, called Mesh-VQ-VAE, provides a low-dimensional discrete latent representation preserving the spatial structure of the mesh. We then propose a transformer-based encoder-decoder model, called VQ-HPS, for learning to solve the HPSE problem using the cross-entropy loss. Once the mesh discrete representation is predicted, we can decode it using the Mesh-VQ-VAE decoder and obtain a full mesh following the SMPL mesh topology [40]. Since the Mesh-VQ-VAE is learned on a large human motion capture database [43], it automatically learns to decode smooth and realistic human meshes.
VQ-HPS achieves state-of-the-art performance on the challenging 3DPW [63] benchmark. It outperforms other methods quantitatively while producing qualitative results as realistic as parametric methods (see Fig. 1).
@@ 2.1 Parametric Approaches
Several methods are dedicated to recovering the parameters of a parametric human model, such as SMPL [40]. Optimization techniques iteratively estimate the parameters of a body model based on images or videos, ensuring that the projection of predictions aligns with a set of 2D cues, including 2D skeletons [48, 4, 16, 24], part segmentation [71, 34], or DensePose [20]. Pose and motion priors are commonly incorporated into optimization methods to enhance the realism of predictions [49, 59, 56, 42]. On the contrary, regression methods employ neural networks to predict the parameters of a human body model from input images or videos. Many of these methods leverage convolutional neural networks (CNNs) for extracting image features [31, 29, 25, 30, 73, 36, 68, 8, 26, 69]. Recent works have demonstrated remarkable performance by replacing CNNs with Vision Transformers [13] as seen in [19, 37, 5, 76]. Some methods output probabilistic results, enabling sampling among plausible solutions [54, 33, 2, 17, 55]. While optimization methods typically yield superior results, they come with significantly longer running times than regression methods and require precise initialization and accurate 2D cues. One limitation in training regression models lies in the scarcity of datasets containing RGB data with 3D annotations. Prior works have addressed this challenge by employing synthetic data [61, 6, 3, 54, 47] or pseudo-labels [24, 44, 34] for training their models.
While parametric models offer the capability to estimate reasonable human poses, the model parameter space may not be the most suitable focus for predicting human pose and shape [32, 12]. Recognizing these limitations inherent in parametric approaches has spurred the development of non-parametric methods.
@@ 2.2 Non-parametric Approaches
Several works have explored methods for directly predicting 3D meshes without relying on the parameters of a human body model [32, 12, 45, 39, 38, 7]. In earlier approaches, regression architectures based on Graph Convolutional Neural Networks (GCNNs) were proposed, utilizing a graph structure derived from the topology of the SMPL human mesh [32, 45, 39]. Recent advancements have leveraged Transformer architectures, capitalizing on attention mechanisms to capture relationships between joints and vertices. While approaches like [39, 38] have introduced encoder-based strategies that concatenate image features and mesh tokens for predicting 3D coordinates, FastMETRO [7] presented an encoder-decoder architecture, effectively disentangling image and mesh modalities. Recently, [14] introduced a token pruning strategy to enhance the efficiency of transformer-based HPSE, and [12] achieved state-of-the-art accuracy in body shape estimation through an optimization-based approach relying on per-vertex neural features.
In this work, we introduce a non-parametric approach to Human Pose and Shape Estimation (HPSE). Our objective is to estimate the vertices of a human body mesh, adhering to the SMPL topology [40]. In contrast to all prior works, our method involves predicting the mesh through a discrete latent representation, reframing HPSE as a classification problem. Although exploiting the discriminative power of classification networks has already been proposed for the Human Pose Estimation (HPE) problem (see below), to the best of our knowledge, this has not been done before for the HPSE problem.
@@ 2.3 Quantization of the Human Pose and Shape
A number of works explored the use of quantization for Human Pose Estimation (HPE).  [35] proposed to discretize horizontal and vertical coordinates for 2D HPE. On the other hand, [51, 50] used anchor poses and refined them for solving the 3D HPE problem. [11] proposed a human pose and shape classification method, but the system was trained on only 12 different postures. Some approaches proposed hand shape classification [28], especially for sign-language recognition following the works of [10]. Some works also proposed face shape classification [53] and head pose estimation [21] using a Support Vector Machine.
Similar to ours, recent works in human motion generation [41, 57, 75, 70] used a VQ-VAE [60] for quantizing human motion. The main difference between these models and Mesh-VQ-VAE is that in these works, a single index encodes a sequence of poses, whereas we use several indices to encode a single pose, allowing for higher precision. Also, none of these works encode the 3D mesh: [41] works on the SMPL parameters, and others only encode a 3D skeleton.
@@ 3 Background
Figure 2: VQ-HPS global process for predicting the mesh given an image. The first step is to predict the camera $\hat{\pi}$ and the rotation $\hat{R}$ from the image $I$. Then, we use the image, the predicted rotation, and the camera to predict the vertices $\hat{V}_{c}$ of the canonical mesh. Finally, $\hat{V}_{c}$ is rotated according to $\hat{R}$ to obtain the final mesh vertices $\hat{V}$. Details about the architecture of VQ-HPS components are also given.
SMPL model. SMPL [40] is a skinned vertex-based model that maps the body shape parameter $\beta\in\mathbb{R}^{10}$ and the pose parameter $\theta\in\mathbb{R}^{72}$ to 3D vertices through the differentiable function $\mathcal{M}(\beta,\theta)$. It outputs the 3D vertices $V\in\mathbb{R}^{6890\times 3}$ of a registered mesh, and 3D joints $J\in\mathbb{R}^{24\times 3}$ can be extracted from the mesh using the joint regressor matrix $\mathcal{J}_{smpl}$. In this work, we do not predict the parameters of the SMPL model, as prior works [32, 9, 12] showed that they are not a suitable target for regression models. However, the mesh predicted by the proposed VQ-HPS model follows the SMPL mesh topology. It allows us to use tools like joint regressors and provides a fair comparison with existing approaches.
Fully convolutional mesh autoencoder. The fully convolutional mesh autoencoder [78] is an autoencoder specifically tailored for handling arbitrary registered mesh data. It relies on the definition of novel convolution and pooling operators with globally shared weights and locally varying coefficients depending on the mesh structure. These variable coefficients play a pivotal role in capturing intricate details inherent to irregular mesh connections, contributing to the model’s exceptional performance in mesh reconstruction. One of the main advantages of the fully convolutional architecture is that the latent codes are localized, which gives a latent space preserving the spatial structure of the mesh. The latent representation of the fully convolutional mesh autoencoder lies in $\mathbb{R}^{N\times d}$ where $N$ is the number of latent vectors, and $d$ is the dimension of latent vectors.
Vector quantized-variational autoencoder. The VQ-VAE [60] is an encoder-decoder model with a discretized latent space. The idea is to learn jointly an encoder, a dictionary of latent codes, and a decoder. The encoder maps the input data $x$ into a latent variable $z\in\mathbb{R}^{N\times d}$. We then discretize $z$ using a learned dictionary of $S$ latent codes. We can then write $z_{d}\in\mathbb{R}^{N\times d}$ where each vector of $z$ is replaced by the closest latent code, or $z_{q}\in\{1,\dots,S\}^{N}$, where the index of the closest latent code replaces each vector of $z$. The decoder reconstructs $x$ from the discrete latent representation $z_{d}$ and the learned dictionary.
@@ 4 Method
Figure 3: figure
Mesh-VQ-VAE reconstruction error. Samples of reconstruction on the 3DPW test set. The error is in cm and corresponds to the Euclidean distance between the original mesh and the corresponding vertex in the reconstruction.
Figure 3: figure
@@ 4.1 Proposed HPSE method
We propose a novel classification-based method for HPSE. Our goal is to predict an oriented 3D mesh from an image. VQ-HPS consists of an encoder-decoder architecture, predicting the human mesh discrete representation of the introduced Mesh-VQ-VAE from image features. We believe that this is the most adapted architecture for predicting our discrete latent representation, with encoder tokens corresponding to image patches and the decoder tokens corresponding to parts of the body and indices in the latent space. In order to ease the low-dimensional representation learning of the mesh, the predicted mesh is non-oriented and centered on the origin (see Sec. 4.2): we call it a canonical mesh. To obtain the final oriented mesh, we then need to predict the rotation $R\in\mathbb{R}^{3\times 3}$, and for better alignment with the image, we also regress the perspective camera $\pi=[s,t]\in\mathbb{R}^{3}$ where $s$ is a scale parameter and $t$ is a 2D translation. The overall method is shown in Fig. 2, and we will now proceed to detail each of its primary components.
@@
Mesh-VQ-VAE. For learning discrete representations of meshes, we build on the fully convolutional mesh autoencoder [78] (see Sec. 3) for encoding the full mesh vertices $V\in\mathbb{R}^{6890\times 3}$ to a latent representation $z\in\mathbb{R}^{N\times d}$. We add a quantization step in the latent space similar to [60] (see Sec. 3). $z$ is mapped to the discrete latent representation $z_{q}\in\{1,\dots,S\}^{N}$. While the fully convolutional architecture preserves the spatial structure of the mesh, the added quantization step allows us to view the HPSE as a classification task. Our Mesh-VQ-VAE Fig. 2 can be seen as a VQ-VAE [60] where the encoder and decoder networks are the fully convolutional mesh autoencoder.
@@
Feature extractors. The first step for image-based HPSE is to extract features from the image. We use CNN backbones to preserve the spatial structure of the image, and we obtain features $X\in\mathbb{R}^{H\times W\times C}$, where $C$ is the number of channels of the backbone and $H\times W$ is the spatial dimension size. We use two backbones: one for predicting the global rotation and another for the pose and shape prediction.
@@
Rotation and camera prediction. We start by predicting the mesh rotation and the perspective camera parameters (see again Fig. 2). These predictions depend on the image features and on an initial body pose $p\in\mathbb{R}^{17\times 3}$ following the Human3.6M [23] joints layout. We predict the rotation $\hat{R}$ and the weak perspective camera parameters $\hat{\pi}$.
@@
Latent canonical mesh regressor encoder. The transformer encoder inputs are the features extracted by the CNN backbone. Before being fed to the Transformer encoder, we apply a 1D-convolution on the image features to make them of dimension $X^{\prime}_{mesh}\in\mathbb{R}^{H\times W\times D}$ where $D$ is the hidden state size of the Transformer. These features are flattened to obtain $HW$ tokens of dimension $D$, and then we add positional encoding. The obtained tokens are fed to a Transformer encoder, using self-attention between all image tokens to output encoded image features $X_{feat}\in\mathbb{R}^{HW\times D}$.
@@
Latent canonical mesh regressor decoder. The transformer decoder takes as inputs $N$ learned mesh tokens $M_{t}$ of size $D$, each responsible for predicting a dimension of the Mesh-VQ-VAE discrete latent representation. The Transformer decoder consists of self-attention between learned tokens and cross-attention with image features. It outputs latent mesh features $z_{feat}\in\mathbb{R}^{N\times D}$. Then (see Fig. 2), to obtain the logits $\hat{z}_{log}\in\mathbb{R}^{N\times S}$, we rely on the mesh features as well as on the previously predicted rotation and camera. We obtain the predicted discrete representation $\hat{z}_{q}\in\{1,\dots,S\}^{N}$ by applying an $\arg\max(\cdot)$ operation.
@@
Reconstructing the full mesh. From the mesh discrete latent representation $\hat{z}_{q}\in\{1,\dots,S\}^{N}$, we use the decoder of the introduced Mesh-VQ-VAE to reconstruct the vertices of a full canonical mesh $\hat{V}_{c}\in\mathbb{R}^{6890\times 3}$. To obtain the vertices $\hat{V}$ of the oriented mesh in the frame coordinates, we apply the predicted rotation $\hat{R}$ to the vertices. This process is shown in Fig. 2.
4.2 Implementation details
Mesh-VQ-VAE. The Mesh-VQ-VAE architecture (see Fig. 2) is adapted from the fully convolutional mesh autoencoder of [78]. This model encodes a mesh to a latent representation $z\in\mathbb{R}^{N\times d}$ with $N=54$ and $d=9$. The quantization step is performed with a dictionary of size $S=512$.
Feature extractors. In order to provide a fair comparison with previous methods, our CNN backbones are ResNet-50 [22] pre-trained on ImageNet [52]. For the canonical mesh prediction, we remove the last fully connected layer to obtain features $X_{mesh}\in\mathbb{R}^{H\times W\times C}$ with $H=W=7$, $C=2048$. For the rotation and camera prediction, we keep the full Resnet-50 to obtain features $X_{rot}\in\mathbb{R}^{C}$.
Rotation and camera predictors. Inspired by [25], the network for predicting the rotation and the mesh consists of a multilayer perceptron (MLP) regression module composed of two fully connected layers with 1024 neurons following the CNN backbone (see Fig. 2). The image feature $X_{rot}\in\mathbb{R}^{C}$ is concatenated with the flattened pose $p\in\mathbb{R}^{17*3}$ before being fed to the MLP. The rotation is predicted in the 6d-rotation format [77], and $p$ is initialized with the SMPL T-pose.
Latent canonical mesh regressor. The Transformer’s hidden dimension is $D=512$, and all multi-layer-perceptrons in the encoder and decoder have a hidden size of 1024. We use sinusoidal positional encoding for the input tokens.

### v2

# suppSupplementary References

@@ Abstract
Previous works on Human Pose and Shape Estimation (HPSE) from RGB images can be broadly categorized into two main groups: parametric and non-parametric approaches. Parametric techniques leverage a low-dimensional statistical body model for realistic results, whereas recent non-parametric methods achieve higher precision by directly regressing the 3D coordinates of the human body mesh. This work introduces a novel paradigm to address the HPSE problem, involving a low-dimensional discrete latent representation of the human mesh and framing HPSE as a classification task. Instead of predicting body model parameters or 3D vertex coordinates, we focus on predicting the proposed discrete latent representation, which can be decoded into a registered human mesh. This innovative paradigm offers two key advantages. Firstly, predicting a low-dimensional discrete representation confines our predictions to the space of anthropomorphic poses and shapes even when little training data is available. Secondly, by framing the problem as a classification task, we can harness the discriminative power inherent in neural networks. The proposed model, VQ-HPS, predicts the discrete latent representation of the mesh. The experimental results demonstrate that VQ-HPS outperforms the current state-of-the-art non-parametric approaches while yielding results as realistic as those produced by parametric methods when trained with little data. VQ-HPS also shows promising results when training on large-scale datasets, highlighting the significant potential of the classification approach for HPSE. See the project page at https://g-fiche.github.io/research-pages/vqhps/
Figure 1: VQ-HPS formulates the human pose and shape estimation problem as a classification task in a vector-quantized latent space. We present the results of VQ-HPS on two challenging scenarios with in-the-wild conditions and poor illumination, comparing its performance to that of HMR [1], Cliff [2] and FastMETRO [3] when trained on little data.
@@ 1 Introduction
To alleviate these issues, several works switched to methods inspired by 3D pose estimation models that predict 3D coordinates directly. Earlier methods predicted the 6890 vertices of the full SMPL [9] mesh using graph convolutional neural networks (GCNNs) modeling the mesh structure and focusing on local interaction between neighboring vertices [20, 19]. While [24] used Transformers [25] to model global interactions between joints and vertices, others argued that a hybrid architecture mixing Graph Convolutional Neural Networks (GCNNs) and Transformers would enable modeling both local and global interactions [23]. More recently, FastMETRO [3] proposed a Transformer-based encoder-decoder architecture to disentangle image and mesh features and to predict 3D coordinates of body joints and a coarse mesh that can be upsampled to the full SMPL body mesh. Significantly different from prior works, LVD [18] proposed an optimization-based approach estimating each vertex position independently by predicting vertex displacement with neural fields. Despite proposing alternatives to model-based approaches, these methods also present some drawbacks: 1) Approaches regressing all vertices of the body mesh at once lack global interaction modeling when using GCNNs [23] and have a very high computational cost when using Transformers [3, 26]; 2) Regression-based methods sometimes output noisy meshes, some of them regress the SMPL parameters from the predicted mesh to obtain smoother predictions, but it comes with a loss of accuracy [3, 19, 20]. This problem is even more glaring when little training data is available, with non-anthropomorphic predictions as displayed in Fig. 1 for[3]; 3) Methods regressing 3D vertices are very sensible to the distribution shift between training and test data [24] (see Sec. 5.4); 4) LVD [18] is real-time and obtains state-of-the-art results for shape estimation, but is not adapted to extreme poses.
@@
This work introduces a method significantly different from all prior human pose and shape estimation (HPSE) approaches. Instead of predicting the parameters of a human body model or 3D coordinates, we learn to predict a discrete latent representation of 3D meshes, transforming the HPSE into a classification problem in which we can exploit the originally targeted discriminative power of Transformers, which has been proven unmatched in natural language processing. For learning our discrete latent representation of meshes, we build on the vector quantized-variational autoencoder (VQ-VAE) [27] framework and adapt it to the fully convolutional mesh autoencoder proposed in [28]. The encoder of the proposed model, called Mesh-VQ-VAE, provides a low-dimensional discrete latent representation preserving the spatial structure of the mesh. We then propose a Transformer-based encoder-decoder model, called VQ-HPS, for learning to solve the HPSE problem using the cross-entropy loss. Once the mesh discrete representation is predicted, we can decode it using the pre-trained Mesh-VQ-VAE decoder and obtain a full mesh following the SMPL mesh topology [9]. Since the Mesh-VQ-VAE is pre-trained on a large human motion capture database [29], it automatically learns to decode smooth and realistic human meshes. This is particularly interesting when training with little data: VQ-HPS learns to predict sequences of indices corresponding to realistic meshes early in the training process as demonstrated in Appendix C.
In the context of few training data availability, VQ-HPS achieves state-of-the-art performance on the challenging in-the-wild 3DPW [30] and EMDB [31] benchmarks: it significantly outperforms other methods quantitatively while producing qualitative results as realistic as parametric methods (see Fig. 1). Moreover, it shows promising results when trained on standard large-scale datasets, enhancing the significant potential of the classification-based approach for solving the HPSE problem.
@@
  • Code and trained models will be released upon acceptance.
@@ 2.1 Parametric Approaches
Several methods are dedicated to recovering the parameters of a parametric human model, such as SMPL [9]. Optimization techniques iteratively estimate the parameters of a body model based on images or videos, ensuring that the projection of predictions aligns with a set of 2D cues, including 2D skeletons [12, 14, 32, 33], part segmentation [34, 15], or DensePose [35]. Pose and motion priors are commonly incorporated into optimization methods to enhance the realism of predictions [16, 36, 37, 38]. On the contrary, regression methods employ neural networks to predict the parameters of a human body model from input images or videos. Many of these methods leverage convolutional neural networks (CNNs) for extracting image features [39, 40, 1, 41, 42, 2, 43, 44, 45, 46, 47, 48]. Recent works have demonstrated remarkable performance by replacing CNNs with Vision Transformers [49] as seen in [17, 50, 51, 52]. Some methods output probabilistic results, enabling sampling among plausible solutions [53, 54, 55, 56, 57]. While optimization methods typically yield superior results, they come with significantly longer running times than regression methods and require precise initialization and accurate 2D cues. One limitation in training regression models is the scarcity of RGB data with 3D annotations. Prior works have addressed this challenge by employing synthetic data [58, 59, 60, 53, 61] or pseudo-labels [33, 62, 15] for training their models.
While parametric models can estimate reasonable human poses, the model parameter space may not be the most suitable focus for predicting human pose and shape [19, 18]. Recognizing these limitations inherent in parametric approaches has spurred the development of non-parametric methods.
@@ 2.2 Non-parametric Approaches
Several works have explored methods for directly predicting 3D meshes without relying on the parameters of a human body model [19, 18, 63, 23, 24, 3]. In earlier approaches, regression architectures based on GCNNs were proposed, utilizing a graph structure derived from the topology of the SMPL human mesh [19, 63, 23]. Recent advancements have leveraged Transformer architectures, capitalizing on attention mechanisms to capture relationships between joints and vertices. While approaches like [23, 24] have introduced encoder-based strategies that concatenate image features and mesh tokens for predicting 3D coordinates, FastMETRO [3] presented an encoder-decoder architecture, effectively disentangling image and mesh modalities. Recently, [26] introduced a token pruning strategy to enhance the efficiency of Transformer-based HPSE, and [18] achieved state-of-the-art accuracy in body shape estimation through an optimization-based approach relying on per-vertex neural features.
This work introduces a non-parametric approach to HPSE. Our objective is to estimate the vertices of a human body mesh, adhering to the SMPL topology [9]. In contrast to all prior works, our method involves predicting the mesh through a discrete latent representation, reframing HPSE as a classification problem. Although exploiting the discriminative power of classification networks has already been proposed for the Human Pose Estimation (HPE) problem (see Sec. 2.3), to our knowledge, this has not been done before for the HPSE problem.
@@ 2.3 Quantization of the Human Pose and Shape
Some works explored quantization for HPE.  [64] proposed to discretize horizontal and vertical coordinates for 2D HPE. On the other hand, [65, 66] used anchor poses and refined them for solving the 3D HPE problem. [67] proposed a human pose and shape classification method, but the system was trained on only 12 different postures. Some approaches proposed hand shape classification [68], especially for sign-language recognition following the works of [69]. Some works also proposed face shape classification [70] and head pose estimation [71] using a Support Vector Machine.
Recent works in human motion generation [72, 73, 74, 75] used a VQ-VAE [27] for quantizing human motion. The main difference with the proposed Mesh-VQ-VAE is that a single index encodes a sequence of poses in these works. In contrast, we use several indices to encode a single pose, allowing for higher precision. Also, none of these works encode the 3D mesh: [72] uses the SMPL parameters, and others only encode a 3D skeleton. Furthermore, human motion forecasting and generation tasks differ significantly from HPSE.
Concurrently with the present work, TokenHMR [76] proposes to quantize human pose for HPSE. As in our case, this tokenization acts as a pose prior, using a dictionary of valid pose tokens. However, the tokenization of TokenHMR differs from ours as it happens on the SMPL [9] pose parameter while we quantize the 3D mesh. Another major difference between VQ-HPS and TokenHMR is how the tokenized pose is used. TokenHMR uses it as an intermediate representation while still solving the HPSE problem as a regression task: similar to prior works in HPSE, the training targets are the SMPL pose and shape parameters and the vertices’ coordinates. We propose to frame HPSE as a classification task: the unique training target for the mesh is its discrete quantized representation. Experiments show that VQ-HPS achieves better results than TokenHMR using less training data and a less powerful backbone.
@@ 3 Background
SMPL model. SMPL [9] is a skinned vertex-based human body model that maps the body shape parameter $\beta\in\mathbb{R}^{10}$ and the pose parameter $\theta\in\mathbb{R}^{72}$ to 3D vertices through the differentiable function $\mathcal{M}(\beta,\theta)$. It outputs the 3D vertices $V\in\mathbb{R}^{6890\times 3}$ of a registered mesh, and 3D joints $J\in\mathbb{R}^{24\times 3}$ can be extracted from the mesh using the joint regressor matrix $\mathcal{J}_{smpl}$. In this work, we do not predict the parameters of the SMPL model, as prior works [19, 20, 18] showed that they are not a suitable target for regression models. However, the mesh predicted by the proposed VQ-HPS model follows the SMPL mesh topology. It allows us to use tools like joint regressors and provides a fair comparison with existing approaches.
Fully convolutional mesh autoencoder. The fully convolutional mesh autoencoder [28] is an autoencoder specifically tailored for handling arbitrary registered mesh data. It relies on the definition of novel convolution and pooling operators with globally shared weights and locally varying coefficients depending on the mesh structure. These variable coefficients are pivotal in capturing intricate details inherent to irregular mesh connections, contributing to the model’s performance in mesh reconstruction. One of the main advantages of fully convolutional architecture is that the latent codes are localized, which gives a latent space preserving the spatial structure of the mesh. The latent representation of the fully convolutional mesh autoencoder lies in $\mathbb{R}^{N\times L}$ where $N$ is the number of latent vectors, and $L$ is the dimension of latent vectors.
Vector quantized-variational autoencoder. The VQ-VAE [27] is an encoder-decoder model with a discretized latent space. The idea is to learn jointly an encoder, a dictionary of latent codes, and a decoder. The encoder maps the input data $x$ into a latent variable $z\in\mathbb{R}^{N\times L}$. We then discretize $z$ using a learned dictionary of $S$ latent codes of dimension $L$. We can then write $z_{d}\in\mathbb{R}^{N\times L}$ where each vector of $z$ is replaced by the closest latent code, or $z_{q}\in\{1,\dots,S\}^{N}$, where the index of the closest latent code replaces each vector of $z$. The decoder reconstructs $x$ from the discrete latent representation $z_{d}$ and the learned dictionary.
@@ 4 Method
Figure 2: VQ-HPS global process for predicting the mesh given an image. We first predict the camera $\hat{\pi}$ and the rotation $\hat{R}$ from the image $I$. Then, we use the image, the predicted rotation, and the camera to predict the vertices $\hat{V}_{c}$ of the canonical mesh. Finally, $\hat{V}_{c}$ is rotated according to $\hat{R}$ to obtain the final mesh vertices $\hat{V}$.
@@ 4.1 Proposed HPSE method
We propose a novel classification-based method for HPSE. Our goal is to predict an oriented 3D mesh from an image. VQ-HPS consists of an encoder-decoder architecture, predicting the human mesh discrete representation of the introduced Mesh-VQ-VAE from image features. We believe this is the most adapted architecture for predicting our discrete latent representation, with encoder tokens corresponding to image patches and decoder tokens corresponding to body parts and indices in the latent space. To ease the low-dimensional representation learning of the mesh, the predicted mesh is non-oriented and centered on the origin (see Appendix A): we call it a canonical mesh. To obtain the final oriented mesh, we then need to predict the rotation $R\in\mathbb{R}^{3\times 3}$, and for better alignment with the image, we also regress the perspective camera $\pi=[s,t]\in\mathbb{R}^{3}$ where $s$ is a scale parameter and $t$ is a 2D translation. The overall method is shown in Fig. 2, and we will now detail each of its primary components.
@@
Mesh-VQ-VAE. For learning discrete representations of meshes, we build on the fully convolutional mesh autoencoder [28] (see Sec. 3) for encoding the full canonical mesh vertices $V_{c}\in\mathbb{R}^{6890\times 3}$ to a latent representation $z\in\mathbb{R}^{N\times L}$. We add a vector quantization step in the latent space similar to [27] (see Sec. 3), which maps $z$ to the discrete latent representation $z_{q}\in\{1,\dots,S\}^{N}$. While the fully convolutional architecture preserves the spatial structure of the mesh, the added quantization step allows us to view the HPSE as a classification task as we aim to predict the indices of the latent mesh representation given an image. Our Mesh-VQ-VAE Fig. 2 can be seen as a VQ-VAE [27] whose architecture corresponds to the fully convolutional mesh autoencoder.
@@
Feature extractors. The first step for image-based HPSE is to extract features from the image. We use CNN backbones to preserve the spatial structure of the image, and we obtain features $X\in\mathbb{R}^{H\times W\times C}$, where $C$ is the number of channels of the backbone and $H$ and $W$ are the spatial dimension sizes. We use two feature extractors. The feature extractor of the camera and rotation predictors gives $X_{r}ot$. The feature extractor gives $X_{mesh}$ with $W=1$ in the latent canonical mesh predictor.
@@
Rotation and camera prediction. We start by predicting the mesh rotation and the perspective camera parameters (see again Fig. 2). These predictions depend on the image features and an initial body pose $p\in\mathbb{R}^{17\times 3}$ following the Human3.6M [77] joints layout. We predict the rotation $\hat{R}$ and the weak perspective camera parameters $\hat{\pi}$.
@@
Latent canonical mesh regressor encoder. We then predict the discrete latent representation of the canonical mesh. The Transformer encoder inputs are the features extracted by the CNN backbone. Before being fed to the Transformer encoder, we apply a 1x1 convolution on the image features to make them of dimension $X^{\prime}_{mesh}\in\mathbb{R}^{H\times W\times D}$ where $D$ is the hidden state size of the Transformer. These features are flattened to obtain $HW$ tokens of dimension $D$, and then we add positional encoding. The obtained tokens are fed to a Transformer encoder, using self-attention between all image tokens to output encoded image features $X_{feat}\in\mathbb{R}^{HW\times D}$.
@@
Latent canonical mesh regressor decoder. The Transformer decoder takes as inputs $N$ learned mesh tokens $M_{t}$ of size $D$, each responsible for predicting an index of the Mesh-VQ-VAE discrete latent representation. We need to solve $N$ classification problems, one for each index. Each problem is a $S$-class classification, $S$ corresponding to the size of the Mesh-VQ-VAE codebook. The Transformer decoder consists of self-attention between learned tokens and cross-attention with image features. It outputs latent mesh features $z_{feat}\in\mathbb{R}^{N\times D}$. Then (see Fig. 2), to obtain the logits $\hat{z}_{log}\in\mathbb{R}^{N\times S}$, we rely on the mesh features as well as on the previously predicted rotation and camera. We obtain the predicted discrete representation $\hat{z}_{q}\in\{1,\dots,S\}^{N}$ by applying an $\arg\max(\cdot)$ operation.
@@
Reconstructing the full mesh. From the discrete latent mesh representation $\hat{z}_{q}\in\{1,\dots,S\}^{N}$, we use the decoder of the introduced Mesh-VQ-VAE to reconstruct the vertices of a full canonical mesh $\hat{V}_{c}\in\mathbb{R}^{6890\times 3}$. We apply the predicted rotation $\hat{R}$ of the oriented mesh in the frame coordinates to the vertices to obtain the vertices $\hat{V}$. This process is shown in Fig. 2.
Figure 3: Mesh-VQ-VAE reconstruction error. Samples of reconstruction on the 3DPW test set. The error is in cm and corresponds to the Euclidean distance between the reconstruction’s original mesh and the corresponding vertex.